OpenStudy (anonymous):

Which answer shows this equation written in standard form? 9 – 4(x – y) = –6x + 6 A. –10x + 4y = –3 B. 2x – 4y = 3 C. 2x + 4y = 3 D. 2x + 4y = –3

OpenStudy (jordanloveangel):

B 9-4x+4y = -6x+6 +6x +6x ______________ 9+2x+4y = 6 -9 -9 _______________ 2x+4y = -3

OpenStudy (anonymous):

The correct answer is, B. I worked it out on paper for you.

OpenStudy (anonymous):

Thanks:)